The answers circled are wrong show the work to arrive to one of the choices please...


media%2F2c4%2F2c43ca0d-5b71-4238-ba07-77 The answers circled are wrong show the work to arrive to one of the choices please will rate more points.
Answer #1

1)
P(A) = 0.35

P(B|A) = 3 P(B|A')
P(B and A) /P(A) = 3 P(B and A')/P(A')
P(B and A) /P(A) = 3 (P(B) - P(A and B))/(1- P(A))
P(A and B) = 3 * 0.35/0.65 ( P(B) - P(A and B))
P(A and B) (1 + 1.615384) = 1.615384 P(B)
P(A and B) /P(B) = 1.615384/2.615384 = 0.61764

P(A|B) = P(A and B)/P(B)
=0.61764
=0.62

Please rate
Please post next question again